a new cruise ship line has just launched 3 new ships: the pacific paradise, the Caribbean paradise, and the Mediterranean paradise. The Caribbean paradise has 37 more deluxe staterooms than the pacific paradise. The Mediterranean paradise has 38 fewer deluxe staterooms than four times the number of deluxe staterooms on the pacific paradise. find the number of the deluxe staterooms for each of the ships if the total number of deluxe staterooms for all three ships is845
Answer by Theo(13342)

p = number of deluxe staterooms on the pacific paradise
c = number of deluxe staterooms on the caribbean paradise
m = number of deluxe staterooms on the mediterranean paradise.

c = p + 37
m = 4p - 38
p + c + m = 845
substitute for c and m to get:
p + p + 37 + 4p - 38 = 845
combine like terms to get:
6p - 1 = 845
add 1 to both sides of the equation to get:
6p = 846
divide both sides of the equation by 6 to get:
p = 846 / 6 = 141
c = p + 37 = 141 + 37 = 178
m = 4p - 38 = 4*141 - 38 = 526
p + c + m = 845 confirming the solutions are good.